# Task Diversity Shortens the ICL Plateau

### v1

@@ 4.2.2 Generality of our hypothesis
$L(W,V)=\sum_{m=1}^{k}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\left(v_{i}^{\intercal}\sigma(Wx)-u_{i}^{\intercal}\sigma(A_{\star}x)\right)^{2}\right],\quad U=\left[\begin{matrix}u_{1}&\cdots&u_{k}\end{matrix}\right],\,V=\left[\begin{matrix}v_{1}&\cdots&v_{k}\end{matrix}\right].$
@@
Figure 7: (Left) Language ICL task. Previous work Fu et al. (2024) identified the difficulty of learning the WordSelection$(4)$ task. We found that mixing it with WordLength or WordSelection$(2)$ reduces the plateau. Refer to Appendix C.1 for further details. (Right) Feature learning setup. For the toy model described in Section 4.2.2, multi-task feature learning significantly shortens the loss plateau. Refer to Appendix E for further details.
@@ 4.2.3 The common structure is not just an induction head
Note that the induction head is precisely the mechanism for solving this Retrieval task. We conduct a checkpoint experiment for the Retrieval tasks, allowing the models to learn the induction head. To train continuous ICL tasks, we use a checkpoint model pre-trained with Gaussian Retrieval. To train Boolean ICL tasks, we use a checkpoint model pre-trained with Boolean Retrieval. We find that the checkpoint does not significantly shorten the ICL tasks’ plateaus. This result demonstrates that the common structure shared by other ICL tasks is not just an induction head. (It is unclear whether an induction head is useful for our ICL tasks at all.) For further details, please refer to Appendix D.2.
@@ 5 In-context learning tasks
  • ReLU Regression. =$\mathcal{F}=\{f\,|\,f(x)=\mathrm{ReLU\,}(w^{\intercal}x)\}$. $\mathcal{D}_{\mathcal{F}}$: Each elements of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
  • Decision Tree. Consider a full binary tree of fixed depth$=4$. The values of leaf nodes and branch nodes are independently sampled from $\mathcal{N}(\mu,1)$ and $\mathrm{Unif\,}(1,\dots,d)$, respectively. When $x$ traverses through the tree, at each branch node with index $i$, we move right if $x[i]>0$ and move left otherwise. $f(x)$ corresponds to the value of the leaf node reached at the end of the traversal.
@@
  • Conjunction. =$\mathcal{F}=\{f\,|\,f(x)=\left(\wedge_{i\in A}x[i]\right)\wedge\left(\wedge_{i\in B}\bar{x}[i]\right)\}$. $\mathcal{D}_{\mathcal{F}}$: $A$ is an uniformly sampled subset of $\{1,\dots,d\}$. Thereafter, we uniformly sample a subset $B\subseteq\{1,\dots,d\}\backslash A$. .
  • Disjunction. =$\mathcal{F}=\{f\,|\,f(x)=\left(\vee_{i\in A}x[i]\right)\wedge\left(\wedge_{i\in B}\bar{x}[i]\right)\}$. $\mathcal{D}_{\mathcal{F}}$: $A$ is an uniformly sampled subset of $\{1,\dots,d\}$. Thereafter, we uniformly sample a subset $B\subseteq\{1,\dots,d\}\backslash A$.
@@ Acknowledgement
We thank Wonnjong Rhee for prompting this project as part of a course at Seoul National University. We are also grateful to the participants of the KAIST-SNU workshop for their discussions at the early stages of this work. Special thanks to Chanwoo Park for sharing his empirical insights. We thank Minkyu Kim for providing feedback about figures. We also thank Hyunsik Chae, Junsu Kim, and Jaesung Park for reviewing the manuscript and providing valuable feedback.
@@ Appendix A Experimental details
We use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of $0.0001$. The batch size $B=64$ corresponds to each task. Regarding the choice of $\ell(\cdot,\cdot)$, we use mean-squared error loss for continuous ICL tasks and cross-entropy loss for boolean ICL tasks. The input dimensions considered are $d=10$ and $15$.
@@
Appendix B Additional tables
To demonstrate the generality of our claim, we conducted extensive experiments beyond those shown in Table 1. The following tables summarize results for transformer (input dimension $d=15$), Mamba (input dimension $d=10$), and Hyena (input dimension $d=10$). We report two metrics for each run: time to escape the plateau and the time to complete the learning (written in parentheses).
Boolean Tasks Continuous Tasks
Number of tasks Sparse Parity(2) Sparse Parity(3) Conjunction Disjunction Linear Regression Quadratic Regression ReLU Regression Sparse Linear Regression Decision Tree
1 $>1000$k $>1000$k $46.9$k ($66.6$k) $22.6$k ($50.7$k) $58.8$k ($61.3$k) $25.0$k ($25.1$k) $4.1$k ($5.4$k) $31.6$k ($35.3$k) $32.4$k ($86.2$k)
2 $38.9$k ($52.6$k) $104.7$k ($148.2$k) $7.5$k ($16.5$k) $5.0$k ($11.6$k) $12.4$k ($13.9$k) $14.8$k ($17.2$k) $4.1$k ($5.1$k) $4.8$k ($5.4$k) $26.6$k ($70.4$k)
8 $4.0$k ($4.8$k) $10.1$k ($13.1$k) $2.8$k ($8.5$k) $2.7$k ($8.7$k) $2.3$k ($3.7$k) $4.0$k ($4.3$k) $2.4$k ($3.9$k) $2.2$k ($3.6$k) $13.1$k ($44.4$k)
9 $3.6$k ($4.1$k) $10.9$k ($11.6$k) $2.5$k ($7.9$k) $2.5$k ($7.8$k) $2.0$k ($3.2$k) $3.5$k ($3.6$k) $2.0$k ($3.4$k) $1.9$k ($3.2$k) $13.8$k ($50.0$k)
Table 2: Task diversity shortens the ICL plateau. We train a transformer with various combinations of $9$ different tasks ($d=15$).
Boolean Tasks Continuous Tasks
Number of tasks Sparse Parity(2) Sparse Parity(3) Conjunction Disjunction Linear Regression Quadratic Regression ReLU Regression Sparse Linear Regression
1 $7.5$k ($7.7$k) $194.3$k ($196.5$k) $0.4$k ($0.4$k) $0.4$k ($0.4$k) $0.2$k ($0.4$k) $17.7$k ($37.7$k) $0.2$k ($1.0$k) $0.2$k ($0.4$k)
2 $2.9$k ($3.2$k) $14.4$k ($15.6$k) $0.4$k ($0.5$k) $0.3$k ($0.4$k) $0.2$k ($0.5$k) $7.7$k ($22.8$k) $0.3$k ($0.9$k) $0.2$k ($0.5$k)
4 $1.5$k ($1.7$k) $6.0$k ($6.7$k) $0.3$k ($0.4$k) $0.3$k ($0.4$k) $0.2$k ($0.6$k) $3.4$k ($17.1$k) $0.3$k ($0.9$k) $0.2$k ($0.6$k)
8 $2.4$k ($2.4$k) $3.7$k ($4.3$k) $0.3$k ($0.4$k) $0.3$k ($0.4$k) $0.3$k ($0.7$k) $2.8$k ($20.5$k) $0.3$k ($0.9$k) $0.3$k ($0.6$k)
Table 3: Task diversity shortens the ICL plateau (Mamba). We train a Mamba with various combinations of $9$ different tasks ($d=10$).
Boolean Tasks Continuous Tasks
Number of tasks Sparse Parity(2) Sparse Parity(3) Conjunction Disjunction Linear Regression Quadratic Regression ReLU Regression Sparse Linear Regression
1 $1.1$k ($1.3$k) $53.9$k ($58.6$k) $0.2$k ($0.3$k) $0.2$k ($0.3$k) $0.1$k ($0.4$k) $11.1$k ($35.4$k) $0.2$k ($0.7$k) $0.1$k ($0.4$k)
2 $2.0$k ($2.4$k) $23.7$k ($27.2$k) $0.2$k ($0.3$k) $0.2$k ($0.3$k) $0.1$k ($0.4$k) $7.6$k ($33.6$k) $0.2$k ($0.7$k) $0.1$k ($0.4$k)
4 $2.4$k ($2.5$k) $16.8$k ($17.5$k) $0.2$k ($0.3$k) $0.2$k ($0.3$k) $0.1$k ($0.5$k) $6.3$k ($41.7$k) $0.2$k ($0.8$k) $0.1$k ($0.5$k)
8 $3.8$k ($4.0$k) $13.6$k ($14.2$k) $0.2$k ($0.3$k) $0.2$k ($0.3$k) $0.2$k ($0.5$k) $7.1$k ($55.0$k) $0.2$k ($0.9$k) $0.2$k ($0.5$k)
Table 4: Task diversity shortens the ICL plateau (Hyena). We train a Hyena with various combinations of $9$ different tasks ($d=10$).
For Mamba and Hyena, loss plateaus are less consistently observed. Specifically, training on Conjunciton, Disjunction, Linear regression, ReLU Regression, and Sparse Linear Regression rarely exhibit a loss plateau. This tendency, however, does not violate our claim, as we focus on how task diversity shortens the plateau when it does occur.
Appendix C Language ICL tasks
C.1 Natural Language ICL tasks
@@
C.2 Formal Language ICL task: Regbench
@@
Appendix D Checkpoint experiment
D.1 Experimental details
@@
Conjunction Disjunction Boolean Retrieval Linear Regression Quadratic Regression ReLU Regression Sparse Linear Regression Gaussian Retrieval
Iteration number $2.2$k $2.6$k $105$k $19$k $11.7$k $2.8$k $7$k $50$k
@@
After saving the baseline models, we train each task (Conjunction, Disjunction, Linear Regression, Quadratic Regression, ReLU Regression, Sparse Linear Regression, Sparse Parity(2), and Sparse Parity(3)) according to the configuration in Appendix A. To calculate the ratio, we divide the result by the plateau escape time summarized in the first row of Table 1.
@@
D.2 Experiment on Retrieval task
@@
Appendix E Feature learning experiment
@@
Appendix F No-context learning
F.1 Optimal no-context functions
@@
For tasks other than ReLU Regression, we can obtain the closed forms of $g_{\mathcal{F}}^{\star}(x)$ using the above property, as listed follows.
@@
For ReLU Regression, we can estimate $g^{\star}_{\mathcal{F}}(x)$ with small error through a sufficient number of samples from $\mathcal{D}_{\mathcal{F}}$.
F.2 Experimental details and additional figures
@@
for boolean ICL tasks. As shown by the following figures, these measurements are close to 00, which indicates that the model’s output corresponds to task-wise optimal no-context function. We follow Appendix A for other configurations. Moreover, to prevent the optimal no-context be merely zero function, we use $\mu\neq 0$ to parameterize the in-context function classes.

### v2

@@ 4.2.2 Generality of our hypothesis
$L(W,V)$ $=\sum_{m=1}^{k}\mathbb{E}_{x\sim\mathcal{D}_{x}}\left[\left(v_{i}^{\intercal}\sigma(Wx)-u_{i}^{\intercal}\sigma(A_{\star}x)\right)^{2}\right],$
$U$ $=\left[\begin{matrix}u_{1}&\cdots&u_{k}\end{matrix}\right],\,V=\left[\begin{matrix}v_{1}&\cdots&v_{k}\end{matrix}\right].$
@@
Figure 7: (Left) Language ICL task. Previous work Fu et al. (2024) identified the difficulty of learning the WordSelection$(4)$ task. We found that mixing it with WordLength or WordSelection$(2)$ reduces the plateau. Refer to Appendix B.1 for further details. (Right) Feature learning setup. For the toy model described in Section 4.2.2, multi-task feature learning significantly shortens the loss plateau. Refer to Appendix D for further details.
@@ 4.2.3 The common structure is not just an induction head
Note that the induction head is precisely the mechanism for solving this Retrieval task. We conduct a checkpoint experiment for the Retrieval tasks, allowing the models to learn the induction head. To train continuous ICL tasks, we use a checkpoint model pre-trained with Gaussian Retrieval. To train Boolean ICL tasks, we use a checkpoint model pre-trained with Boolean Retrieval. We find that the checkpoint does not significantly shorten the ICL tasks’ plateaus. This result demonstrates that the common structure shared by other ICL tasks is not just an induction head. (It is unclear whether an induction head is useful for our ICL tasks at all.) For further details, please refer to Appendix C.2.
@@ 5 In-context learning tasks
  • LeakyReLU Regression. =$\mathcal{F}=\{f\,|\,f(x)=\mathrm{LeakyReLU\,}(w^{\intercal}x)\}$ with negative slope $\alpha=0.5$. $\mathcal{D}_{\mathcal{F}}$: Each elements of $w\in\mathbb{R}^{d}$ is independently sampled from $\mathcal{N}(\mu,1)$.
@@ Acknowledgement
We thank Wonjong Rhee for prompting this project as part of a course at Seoul National University. We are also grateful to the participants of the KAIST-SNU workshop for their discussions at the early stages of this work. Special thanks to Chanwoo Park for sharing his empirical insights. We thank Minkyu Kim for providing feedback about the figures. We also thank Hyunsik Chae, Junsu Kim, and Jaesung Park for reviewing the manuscript and providing valuable feedback.
Impact statement
This paper advances the understanding of training dynamics in multi-task training, contributing to the broader field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Experimental details
We use the Adam optimizer (Kingma & Ba, 2015) with a learning rate of $0.0001$. We use $B=64$, making each task’s batch size as $64/k$. Regarding the choice of $\ell(\cdot,\cdot)$, we use mean-squared error loss for continuous ICL tasks and cross-entropy loss for boolean ICL tasks. The input dimensions considered are $d=10$ and $15$.
@@
Appendix B Language ICL tasks
B.1 Natural Language ICL tasks
@@
B.2 Formal Language ICL task: Regbench
@@
Appendix C Checkpoint experiment
C.1 Experimental details
@@
Boolean Retrieval Linear Regression Quadratic Regression Sparse Linear Regression Gaussian Retrieval
Iteration number $105$k $19$k $12$k $7$k $50$k
@@
After saving the baseline models, we train each task (Linear Regression, Quadratic Regression, Sparse Linear Regression, Sparse Parity(2), and Sparse Parity(3)) according to the configuration in Appendix A. To calculate the ratio, we divide the result by the plateau escape time summarized in the first row of Table 1.
@@
C.2 Experiment on Retrieval task
@@
Appendix D Feature learning experiment
@@
Appendix E No-context learning
E.1 Optimal no-context functions
@@
For tasks other than LeakyReLU Regression, we can obtain the closed forms of $g_{\mathcal{F}}^{\star}(x)$ using the above property, as listed follows.
@@
For LeakyReLU Regression, we can estimate $g^{\star}_{\mathcal{F}}(x)$ with small error through a sufficient number of samples from $\mathcal{D}_{\mathcal{F}}$.
E.2 Experimental details and additional figures
@@
for boolean ICL tasks. As shown by the following figures, these measurements are close to 00, which indicates that the model’s output corresponds to task-wise optimal no-context function. We follow Appendix A for other configurations. Moreover, to prevent the optimal no-context be merely a zero function, we use $\mu\neq 0$ to parameterize the in-context function classes.